OpenStudy (kitkat1):

Maria wants to solve the following system using the elimination method: 4x + 9y = 59 x + y = 17 What number should the equation x + y = 17 be multiplied by to eliminate y? 9 4 –4 –9

OpenStudy (kitkat1):

Medals!

OpenStudy (kitkat1):

Maria wants to solve the following system using the elimination method: 4x + 9y = 59 x + y = 17 What number should the equation x + y = 17 be multiplied by to eliminate y? 9 4 –4 –9

OpenStudy (anonymous):

Do you know how to use the elimination method?

OpenStudy (kitkat1):

The work of a student to solve a set of equations is shown below: Equation 1: x + 2y = 10 Equation 2: 2x + 5y = 15 Step 1: –2(x + 2y) = –2(10) 2x + 5y = 15 Step 2: –2x – 4y = –10 2x + 5y = 15 Step 3: 1y = 5 Step 4: x + 2(1) = 10 Step 5: x = 8 In which step did the student first make an error? (4 points) Step 3 Step 4 Step 2 Step 1

OpenStudy (kitkat1):

this is the second question can someone explain

OpenStudy (kitkat1):

im not really good at elimination can u show me

OpenStudy (anonymous):

Give me a sec

OpenStudy (kitkat1):

okay

OpenStudy (anonymous):

x+2y=10 2x+5y=15 In order to eliminate a term the need to cancel each other out, I used the x term as it was the easiest. -2*x+2y=10 becomes -2x-4y=-20 So now we have -2x-4y=-20 2x+5y=15 Add those together and what do we get?

OpenStudy (kitkat1):

do i have to substitute?

OpenStudy (anonymous):

Not for elimination method, the point of elimination method is to find the value for a variable by eliminating all other variables. In this case we want to solve for y so we have to get rid of the x on the top and bottom. Since the x on the bottom is positive we have to make the x on top negative so when we add them we get 0x.

OpenStudy (kitkat1):

1y = -5

OpenStudy (kitkat1):

1y = -5 1y 1y Divide = -5

OpenStudy (kitkat1):

so the answer is Step 2 right?

OpenStudy (anonymous):

Almost, we are trying to isolate the y not remove it also in this case you could write it as y=-5 as all variables have an inferred 1 in front of them

OpenStudy (anonymous):

Explain why.

OpenStudy (kitkat1):

because he has –2x – 4y = –10 When the - 10 has to be -20

OpenStudy (anonymous):

Spot On!

OpenStudy (kitkat1):

1y = 5 it should be -5

OpenStudy (kitkat1):

it is asking for the First error

OpenStudy (anonymous):

Yes, indeed it should be y=-5

OpenStudy (kitkat1):

so which one should I choose?

OpenStudy (kitkat1):

it is asking for the first error which is the –2x – 4y = –10 2x + 5y = 15

OpenStudy (kitkat1):

it should be -20

OpenStudy (anonymous):

Correct

OpenStudy (kitkat1):

so the answer is Step 2?

OpenStudy (anonymous):

I believe you just answered your own question :-)

OpenStudy (kitkat1):

Can u help me with last one?

OpenStudy (kitkat1):

Jill wants to solve the following system using the elimination method: y = x + 8 y = 3x – 11 What number should the equation y = x + 8 be multiplied by to eliminate x?

OpenStudy (kitkat1):

I am trying to do it but cant get to an answer

OpenStudy (kitkat1):

–1 3 –3 –2

OpenStudy (anonymous):

What is x in this equation 5-x=0

OpenStudy (kitkat1):

y = x + 8

OpenStudy (kitkat1):

-1?

OpenStudy (anonymous):

What I mean is in 5-x=0 what must x be to make this statement true?

OpenStudy (kitkat1):

5 -5 = 0

OpenStudy (kitkat1):

x =5

OpenStudy (kitkat1):

3-3 = 0 so the answer is 3?

OpenStudy (anonymous):

y = x + 8 y = 3x – 11 Now what must x equal to cancel the positive 3 out?

OpenStudy (kitkat1):

Im not sure

OpenStudy (anonymous):

I think your over thinking this, it's basic addition/subtraction what must happen to have x=0 3+x=0

OpenStudy (kitkat1):

-3 is the answer right?

OpenStudy (anonymous):

Correct

OpenStudy (kitkat1):

okay thank you very much

OpenStudy (anonymous):

No problem, if you need anymore help I would also look on Khan Academy they have resources like this video pretty good stuff helped me get through a lot of my math classes https://www.khanacademy.org/math/cc-eighth-grade-math/cc-8th-systems-topic/cc-8th-systems-elimination/v/addition-elimination-method-1

OpenStudy (anonymous):

Of course don't hesitate to ask either :P

OpenStudy (kitkat1):

I will for sure use that thanks alot